### v1

Abstract
Recent advancements in industrial anomaly detection have been hindered by the lack of realistic datasets that accurately represent real-world conditions. Existing algorithms are often developed and evaluated using idealized datasets, which deviate significantly from real-life scenarios characterized by environmental noise and data corruption such as fluctuating lighting conditions, variable object poses, and unstable camera positions. To address this gap, we introduce the Realistic Anomaly Detection (RAD) dataset, the first multi-view RGB-based anomaly detection dataset specifically collected using a real robot arm, providing unique and realistic data scenarios. RAD comprises 4765 images across 13 categories and 4 defect types, collected from more than 50 viewpoints, providing a comprehensive and realistic benchmark. This multi-viewpoint setup mirrors real-world conditions where anomalies may not be detectable from every perspective. Moreover, by sampling varying numbers of views, the algorithm’s performance can be comprehensively evaluated across different viewpoints. This approach enhances the thoroughness of performance assessment and helps improve the algorithm’s robustness. Besides, to support 3D multi-view reconstruction algorithms, we propose a data augmentation method to improve the accuracy of pose estimation and facilitate the reconstruction of 3D point clouds. We systematically evaluate state-of-the-art RGB-based and point cloud-based models using RAD, identifying limitations and future research directions. The code and dataset could found https://github.com/kaichen-z/PAD.
keywords:
Anomaly Detection, Robotics, Pose Estimation
Figure 1: Illustration of RAD Dataset. Figure (a) illustrates our daily objects family in RAD dataset. Figure (b) demonstrates the dataset collection process employing robot arms.
1 Introduction
The Need for a Real-Life Anomaly Detection Dataset with Robotic Observations: The rapid integration of robotics into modern industrial pipelines has revolutionized manufacturing processes, significantly enhancing efficiency and effectiveness. As the use of robots becomes more widespread, anomaly detection has emerged as a crucial component of these pipelines. However, the development and evaluation of most existing industrial anomaly detection algorithms rely on datasets collected under idealized conditions, where sensors provide perfect observations. Consequently, implementing reliable anomaly detection in real-world robotic systems presents several challenges. One major challenge stems from the presence of noise in robotic observations, which can be attributed to various external and internal factors. External factors, such as fluctuating lighting conditions, varying object poses, and inconsistent backgrounds, can introduce significant noise into the data. Moreover, internal factors, including imperfect camera calibration, suboptimal movement control, and sensor inaccuracies, further complicate the detection process. These sources of noise may compromise the reliability and effectiveness of anomaly detection algorithms.
To address these challenges and bridge the gap between research and real-world applications, there is a pressing need for a benchmark dataset that accurately represents the complexities and variabilities of real-life industrial scenarios. Such a dataset should encompass diverse conditions and account for the various sources of noise that can affect robotic observations. By providing a realistic and comprehensive testbed, this dataset would enable the development and benchmarking of more robust and effective anomaly detection algorithms.
The availability of a representative benchmark dataset is crucial for advancing the field of industrial anomaly detection using robotics. It would allow researchers and practitioners to evaluate and compare the performance of different algorithms under realistic conditions, facilitating the identification of strengths and weaknesses. Furthermore, it would drive the development of novel approaches that can handle the challenges posed by real-world environments, ultimately leading to more reliable and efficient anomaly detection in industrial settings.
Limitation of Current Datasets: Existing RGB-based datasets such as LOCO-AD [1], MVTec AD [2, 3] and MPDD [4] primarily focus on single-view RGB observations of objects. This approach misses the critical 3D structural information of the real world, thereby limiting their applicability.
In response to the limitations inherent in 2D datasets, rencent research has increasingly focused on the development of 3D anomaly detection datasets, but they also have their own constraints. For instance, MVTec 3D-AD [5] only provides single-view point clouds and suffers from low resolution, which affects the quality and accuracy of the anomaly detection. While Eyescandies [6] offers higher-quality data, it is limited to simulated environments and does not include real-world data. To address the shortcomings of MVTec 3D-AD and Eyescandies, Real3D-AD [7] offers point clouds of entire objects. However, obtaining a complete point cloud of an object in real-life conditions is challenging. In practical industrial scenarios, objects are not typically suspended in mid-air and are often placed randomly, complicating the acquisition of comprehensive point cloud data. Instead of directly providing point cloud information, PAD [8] offers multi-view RGB data, which can be considered a 3D representation akin to Neural Radiance Fields (NeRF) [9]. However, PAD mainly focuses on simulated data and includes only a small and simple real-world dataset. These limitations highlight the need for a more robust and comprehensive dataset that can accurately reflect the complexities and variabilities of real-world industrial environments.
Based on the analysis above, we introduce RAD, a multi-view RGB-based anomaly detection dataset designed to test the robustness and scalability of existing anomaly detection algorithms on real-world data. Our dataset includes 13 different objects and captures over 5000 images from multiple viewpoints (around 50s), providing a comprehensive testbed for evaluating algorithm performance. RAD records four distinct types of defects, encompassing both texture-based and geometry-based anomalies for each object. This configuration simulates real-life situations where defects may not always be visible from the current viewpoint, and the number of views can vary across different scenarios. To evaluate the quality and versatility of our dataset, we conducted assessments using two distinct types of algorithms: 2D RGB-based and 3D multi-view RGB-based. While 2D RGB-based algorithms have demonstrated impressive performance in RGB-based anomaly detection tasks, their lack of consideration for 3D pose makes them less preferable compared to 3D multi-view RGB-based algorithms for multi-view anomaly detection datasets [10, 8]. However, given that our dataset was collected using real robots shown in Figure. 1, it intrinsically contains noise related to camera calibration and robot positioning. This noise is critical as it significantly affects the performance of the algorithms. Consequently, the direct application of conventional 3D-based algorithms frequently leads to underperformance. To enable effective use of 3D multi-view RGB-based algorithms, we propose a data preprocessing method that accurately estimates camera poses. Building on more accurate pose estimation, we highlight the potential to construct a 3D point cloud using multi-view RGB observations. Despite these efforts, both multi-view RGB-based and 3D point cloud-based algorithms encounter challenges due to variable lighting conditions and the high-specularity of certain objects. This highlights the need for more robust algorithms in future research.
The main contributions of this dataset are summarized as follow:
  • We introduce the first real-life multi-view anomaly detection dataset with robotic observations, featuring over 10 object categories and both texture-based and geometry-based defects. The dataset effectively captures the noisy environments experienced by real robots, bridging the gap between research and real-world applications.
  • We propose the Noisy-Pose-Based Anomaly Detection (NAD) challenge to evaluate algorithm performance under noisy pose conditions and enhance robustness in real-robot scenarios. The NAD challenge provides a standardized benchmark for assessing the effectiveness of anomaly detection algorithms in handling real-world challenges.
  • We propose advanced data preprocessing methods that address the challenges associated with noisy and inconsistent data. These tools facilitate the development and evaluation of novel algorithms that can leverage the rich information available in multi-view and point cloud data.
  • We evaluate the performance of 2D RGB-based, 3D multi-view based, and 3D point cloud-based algorithms on our dataset, providing a detailed analysis of the results. This evaluation highlights the strengths and weaknesses of each approach and identifies areas for future improvement.
Table 1: Comparison of RAD and existing industrial anomaly detection datasets. RGB, D, PC, and NM stand for RGBimage, Depth, Point Cloud, and Normal Maps, respectively. Surface represents the objects that are transparent or specular. ”R.” refers to the dataset with a real-life environment. ”S.” refers to the dataset with simulation environment.
Datasets Representation #Class # Normal #Abnormal #Shoots
MVTec AD [2] (R.) RGB 15 4,096 1,258 1
MPDD [4] (R.) RGB 6 1,064 282 1
VisA [11] (R.) RGB 12 9,621 1,200 1
MVTec 3D-AD [12] (R.) RGB/D/PC 10 2,904 948 1
Eyecandies [6] (S.) RGB/D/NM 10 13,250 2,250 1
Real3D-AD [7] (R.) RGB/D/PC 12 652 602 $360^{\circ}$
PAD [8] (S.) RGB 20 4,960 4,412 Around 20
PAD [8] (R.) RGB 10 271 490 Around 20
Real-IAD [13] (R.) RGB 30 99,721 51,329 5
RAD (Ours) (R.) RGB 13 2,535 2,230 68
Datasets View Color Surface Train. Pose Test. Pose
MVTec AD [2] (R.) Single Diverse Specular 2D 2D
MPDD [4] (R.) Single Diverse Specular 2D 2D
VisA [11] (R.) Single Diverse Specular 2D 2D
MVTec 3D-AD [12] (R.) Single Diverse Specular 3D 2D
Eyecandies [6] (S.) Single Diverse Transparency 3D 2D
Real3D-AD [7] (R.) Multi Diverse Transparency 3D 2D
PAD [8] (S.) Multi Diverse Specular 3D 2D
PAD [8] (R.) Multi Diverse Specular 3D 2D
Real-IAD [13] (R.) Multi Diverse Specular 3D 3D
RAD (Ours) (R.) Multi Diverse Specular 3D 3D
2 Related Work
2.1 Existing Datasets for 3D Anomaly Detection
Anomaly detection and localization is an important task in industrial quality control [14], and there is a wide range of datasets for 2D industrial anomaly detection, such as MVTec AD [2], MPDD [4] and VisA [11]. These datasets spawned many related works on 2D industrial anomaly detection [15, 16, 17], these methods are mainly categorised into image reconstruction [18, 19, 20, 21, 22, 23] and feature embedding[24, 25, 26, 27, 28, 29]. Additionally, the domain of few-shot learning [30, 31] explores enhancements in model performance under fixed data scenarios, whereas anomaly synthesis [32, 33, 34] investigates techniques for generating synthetic anomalies to augment training datasets, thereby improving model robustness. However, since RGB images do not contain spatial information and have a single-view, 2D anomaly detection makes it difficult to detect occlusions or geometric anomalies due to lighting or viewing angle.
Figure 2: Examples of missing anomaly and annotation for each object. The blue frame indicates the normal images in the training dataset. The red frame denotes the anomalous images and ground truth in the testing dataset. The green color is the annotation of the missing region. Each image is a randomly selected single view from each object in RAD.
Some 3D datasets have been proposed to extend anomaly detection from 2D to 3D, which have been compared in Table. 1. MVTec 3D-AD [12] dataset is the first comprehensive 3D dataset with 10 different object categories that is inspired by real-world visual inspection scenarios. However, the MVTec 3D-AD can only provide information from a single viewpoint and lacks RGB information from a full range of viewpoints. In addition, the MVTec 3D-AD dataset has a single background and ambient light, making it difficult to simulate realistic and complex industrial scenes. MPDD [4] dataset supplies multi-view information on metal parts and simulates 12 different types of defects, however, it is constrained in data scale. Eyecandies [6] dataset is a synthetic dataset consisting of synthetic candy views captured under a variety of different lighting conditions. However, it is also limited to single-view and there are differences between the synthetic and real-world data. Real 3D-AD [7] dataset is introduced to expand the object information to 3D space. Real 3D-AD uses a point cloud scanner to scan the object with a full range view and provides the data directly in a point cloud format, thus both the training set and testing set of Real3D-AD contains multi-view information on objects but it only provides spatial information. PAD [8] contains 3D scans of 20 objects, which are used to explore the problem of pose-agnostic anomaly detection. Real-IAD [13] is another latest multi-view datasets that contains 30 objects, but it only provides the information from 5 shooting angles.
Figure 3: Examples of multi-view information. This figure displays a subset of RAD data, illustrating that multiple view observations are provided for each object to support 3D-based anomaly detection algorithms.
2.2 Unsupervised 3D Anomaly Detection
Despite considerable advancements in 2D anomaly detection, the domain of 3D anomaly detection (3D-AD) continues to develop, and substantial challenges persist, particularly in the effective integration of depth and RGB data while maintaining optimal performance. AST [35] combines RGB and depth information to detect defects but the depth information is only used to remove background, anomaly detection is still based on the 2D detection method. M3AD [36] fuses and aligns the feature extracted from the point cloud and RGB images to improve the anomaly detection performance. CPMF [37] combines local geometrical information in 3D modality and global semantic information in the generated pseudo-2D modality to obtain better detection results. Reg3D-AD [7] also employs a dual-feature representation approach, which jointly leverages local features and global representations to detect anomalies in the proposed Real 3D-AD dataset effectively. OmniposeAD [8] addresses pose-agnostic anomaly detection problem by using Neural Radiance Field (NeRF) [9] to reconstruct normal images and refine the pose with image retrieval techniques and iNeRF [38].
2.3 3D Reconstruction
The goal of 3D Reconstruction is to generate a 3D model that accurately captures the shape, geometry, and appearance of a real-world object. The existing 3D reconstruction methods can be categorized as following [39]: Point cloud is a set of 3D points used to represent the surface or structure of an object or scene [40, 41, 42]. On the one hand, sparse point cloud reconstruction methods [43, 44] aim to generate a relatively small number of 3D points that capture the essential structure and geometry of the scene or object. The focus is on identifying and reconstructing key features or landmarks rather than creating a detailed, continuous representation. On the other hand, dense point cloud reconstruction methods [45] aims to create a highly detailed and dense point cloud that captures the complete surface geometry of the object or scene. Voxel-based reconstruction methods represent the surface of a 3D object by using fixed-size cubes as the smallest unit, including 3D voxel grids [46, 47, 48, 49] and octants [50, 51]. A voxel represents a discrete, cubic volume element within a regularly spaced 3D grid and an octree recursively subdivides the 3D space into eight equal-sized octants until a predefined maximum depth or resolution is achieved. Voxel grids are memory-intensive, especially for high-resolution reconstructions. Mesh-based methods define the surface geometry of an object via vertices, edges, and faces [52, 53, 54, 55, 56, 57]. Meshes can provide a high-quality and accurate representation of the object’s surface, capturing fine details and sharp features, while Mesh-based reconstruction is sensitive to noise and incomplete data, and complex surfaces reconstruction may result in large and dense meshes. Neural Radiance Fields (NeRF) [9]-based reconstruction encodes a 3D scene as a continuous function that maps 3D coordinates and viewing directions to color and density values, leveraging a neural network trained on a set of 2D images captured from different viewpoints. This implicit learning allows the network to capture the 3D structure and appearance of the scene. NeRF has been used by PAD [8] to re-rendering high quality images for industrial anomaly detection. NeRF-based 3D surface reconstruction offers several advantages over traditional point cloud-based approaches. NeRF requires only a set of 2D images as input, eliminating the need for explicit 3D scanning or depth information. However, the computational complexity and the requirement for extensive training data and time pose challenges in achieving high-quality reconstruction results.
3 Description of RAD dataset
In this section, we delineate the data acquisition and ground-truth annotation processes for the proposed RAD. Utilizing category and defect information, we will furnish relevant dataset statistics.
3.1 Data Acquisition
Prototype construction
RAD data acquisition process leverages the interactive maneuverability and dynamic data capture capabilities of a robotic arm, facilitating data collection from various angles and perspectives. The robotic arm, equipped with a high-resolution RGB-D camera, captures detailed geometrical and textual data of each object from different angles. Specifically, the prototype construction utilizes a Franka Emika Panda robotic arm as in [58], which carries an ”Intel RealSense Depth camera D415 Webcam 3D” as in [59] with a resolution of 1280$\times$720 pixels, to capture high-resolution images from 68 predefined angles, completing a full $360^{\circ}$ revolution. The object is centrally positioned within the robotic arm’s operational range to ensure consistent imaging across all angles. Each angle’s coordinates are precisely recorded to maintain uniformity in data capture. Nevertheless, owing to the mentioned noises in the introduction, we exclusively offer RGB data and between 40 to 68 views for each category, shown in Figure. 2 and Figure. 3.
In contrast to other data acquisition methods such as using multiple fixed-position cameras or a static point cloud scanner setup, the robotic arm approach offers significant advantages in flexibility and efficiency. Benchmark datasets such as PAD [8] require multiple cameras to cover different angles for shooting, and point cloud scanners [7] need to rotate objects for all-round scanning, while robotic arms can achieve full coverage with a single camera without moving the object. This adaptability makes the robotic arm particularly suitable for industrial assembly lines, where space and setup time are critical considerations. Moreover, automating the data acquisition process with a robotic arm minimizes human error and optimizes the efficiency and reliability of data collection, enhancing the overall robustness of the collection system.
Figure 4: Illustration of Annotation Procedure: This diagram illustrates the annotation process for identifying missing components. Annotations for missing parts primarily rely on comparing defective objects with normal ones.
Details of Data Annotation
To annotate defects, our pipeline comprises two main parts. First, we annotate the defect regions in the abnormal image. Then, we compare this annotation with the normal mask derived from the corresponding normal image. By combining these annotations, we generate a final defect mask. This final mask is overlaid onto the abnormal image to produce the visualization shown in the right part of the Fig. 4. This process allows us to visually highlight and identify defects effectively within the images.
Figure 5: Examples of four defect types. The red frames denote the anomaly regions. Note: entire objects with squeezed anomalies are marked as anomaly regions due to the overall deformation.
Anomaly types
To ensure the dataset’s utility in evaluating robustness against common industrial defects, we simulated four prevalent classes of anomalies as follows:
  • Scratched is generated by scribing the object’s surface with a sharp implement.
  • Missing is simulated by removing segments from the objects.
  • Stained is generated by applying discoloring agents to localized areas of the objects.
  • Squeezed is created by applying pressure to deform parts of the objects.
Figure. 5 demonstrates the example of four different defects of cup and bowl. It is important to note that, due to the varying materials of the items, simulating squeezed anomalies for robust objects presents significant challenges. Consequently, binderclip, box, charger, and phonecase do not include the squeezed defect type.
Table 2: Dataset statistics. The ”Type” category represents the count of distinct object types. ”Single” under ”Attribute” indicates objects typically have one color. ”Multi.” under ”Attribute” indicates objects typically have multiple colors. ”Spec.” under ”Attribute” indicates objects with specular surfaces. ”Miss.” indicates objects with missing defects. ”St.” indicates objects with stained defects. ”Sc.” indicates objects with scratched defects. ”Sq.” indicates objects with squeezed defects.
Category Type Attribute Miss. St. Sc. Sq. Normal Total
Binderclip 2 Single 57 57 57 0 171 342
Bowl 1 Single 55 55 49 55 220 434
Box 1 Single 44 68 68 0 204 384
Can 1 Multi./Spec. 41 41 0 41 164 287
Rubberduck 1 Multi. 68 59 58 68 272 525
Spraybottle 1 Single/Spec. 37 42 38 42 168 327
Cup1 1 Single 44 42 44 44 176 350
Gluebottle 2 Multi./Spec. 32 2 1 44 176 255
Charger 1 Multi./Spec. 28 28 5 0 84 145
Cup2 2 Single 67 28 40 67 268 470
Phonecase 2 Single 68 68 68 68 272 544
Spoon 1 Single 42 42 25 42 168 319
Tennisball 1 Multi. 63 64 64 0 192 383
Labor and time-consuming
The use of a robotic arm for data acquisition represents a significant enhancement in the efficiency and scalability of industrial anomaly detection. By pre-configuring shooting angles within the robotic arm’s control system, we minimize human intervention primarily to the task of replacing objects for capture. This automation not only expedites the data collection process but also ensures consistency and repeatability across diverse datasets, facilitating more extensive and robust data collection endeavors.
Nevertheless, despite the streamlined data acquisition facilitated by the robotic arm, the subsequent phase of data annotation poses substantial challenges in terms of time and resource allocation. In our methodology, the images captured from each configured angle by the robotic arm were annotated using Adobe Photoshop, a process requiring precise manual intervention to identify and label all anomalous pixels. This meticulous process of annotating images from multiple perspectives ensures a comprehensive representation of potential anomalies but took approximately one months to complete for the dataset.
Figure 6: Dataset Metrics. (a) illustrates the pixel-wise ratio within each defect across various categories. (b) shows the total number ratio of each defect in RAD.
3.2 Dataset Statistics
The statistical details of the RAD dataset are outlined in Table. 2 and Figure. 6, which enumerates the dataset categories, the number of training prototypes, and the counts of both normal and various abnormal samples within the test dataset. The RAD dataset comprises a total of 5,848 samples, distributed across 12 distinct objects that encompass multiple domains, including kitchenware, toys, and daily necessities. The categories in RAD dataset exhibit significant diversity in shapes and appearances, providing extensive semantic and geometrical exploration spaces that facilitate a wide range of research objectives. In addition, the data set also includes metadata (camera pose) for each image, which are essential for nuanced analyzes and model training enhancements.
3.3 Ground-truth Annotation
We deliver precise and realistic annotations for each anomaly instance within the test dataset. The annotations are represented as two-dimensional regions, ensuring that the pixel positions in the original image correspond exactly to those in the annotated image. The normal regions of the objects are delineated in white, whereas the anomalous regions are indicated in different colors.
Figure 7: Pipeline of Data Preprocessing. The data preprocessing involves two steps: the first step is pose estimation, followed by 3D reconstruction in the second step.
4 Methodology
Considering the challenges of the NAD problem, directly applying 3D-based anomaly detection methods introduces significant uncertainty. Current multi-view methods, such as MVSNeRF [60], NEUS [61], and 3D Gaussian [62], depend heavily on accurate pose estimation. These multi-view reconstruction methods specifically rely on high-quality pixel-wise matching between different images, which is significantly influenced by the relative positional information between frames. To address this issue, we propose a pipeline to estimate the pose from different views, enabling multi-view based anomaly detection algorithms. Furthermore, we apply 3D reconstruction to create a 3D point cloud, facilitating the use of 3D point cloud-based anomaly detection algorithms. Our pipeline is demonstrated in Figure. 7. It is important to note that our paper primarily focuses on providing an anomaly detection dataset. Our methodology aims to offer a potential solution for future research.
Self-Supervised Pose Estimation: Directly using the poses recorded by the robot may introduce noise and errors due to camera calibration and robot control inaccuracies. To address this issue, we implement the following pose estimation pipeline: Distortion Estimation: Given a series of images { $I\in\mathbb{R}^{H\times W\times 3}$} and corresponding masks $M\in\mathbb{R}^{H\times W}$, we first estimate the distortion parameters based on a radial camera model. Pose Estimation: Using the obtained radial parameters, we perform exhaustive feature matching between image pairs, which helps to the pose estimation. This process helps in reconstructing a sparse point cloud as detailed in [63, 64, 65, 66]. 3D Shape Estimation: With the help of estimated feature and pose information, the sparse point cloud could be estimated. While the sparse point cloud alone is insufficient for evaluating pose estimation quality. Therefore, we apply the patchwise method proposed in [67] to reconstruct a relevant dense point cloud.
By analyzing the reconstructed relevant dense point clouds, we found that directly implementing this pipeline does not always yield reasonable pose estimations when the ground truth pose information is missing. To generate more reasonable pose information in a self-supervised setting, we propose modifying the previous pipeline by adding a preprocessing step and a compactness comparison step. Preprocess: Specifically, we separately apply the following three preprocessing algorithms: brightness adjustment, contrast adjustment, and gamma adjustment. Among them, brightness adjustment can be expressed as:
$I_{\text{adjust}}^{\text{br}}(x,y)=\min\left(\max(I(x,y)+\Delta B,0),255\right),$ (1)
where $\Delta B$ is a constant value.
Contrast adjustment could be written as:
$I_{\text{adjust}}^{\text{con}}(x,y)=\min\left(\max(I(x,y)-M)\cdot C+M,0),255\right),$ (2)
where $M$ is the midpoint or mean intensity of the image. For an 8-bit grayscale image, $M=128$, or you can use the mean intensity of all pixel values as $M$. $C$ is the contrast factor.
Gamma adjustment could be written as:
$I_{\text{adjust}}^{\text{ga}}(x,y)=\min\left(\max(255\cdot\left(\frac{I(x,y)}{255}\right)^{\gamma},255\right),$ (3)
where $\gamma$ is the gamma value that controls the adjustment. The pixel intensities are first normalized by dividing by 255 (for 8-bit images) to map values to the range $[0,1]$, then raised to the power of $\gamma$, and finally scaled back to the range $[0,255]$. Compactness Comparison: By applying different preprocessing algorithms, we obtain different sets of images, including $\{I\}$ and $\{I_{\text{adjust}}^{i}\}$, where $I_{\text{adjust}}^{i}$ indicates images with various adjustment algorithms. Based on these different sets, we can generate distinct sets of pose information. Consequently, different point clouds are estimated. Since we cannot evaluate the pose estimation results directly, we estimate the compactness of the point clouds using the following:
$\text{Compactness}=\frac{\text{Volume}}{\left(\text{Surface\ Area}\right)^{2}},$ (4)
where Volume is the volume of the convex hull of the point cloud. Surface Area is the surface area of the convex hull.
This compactness value allows us to assess the pose estimation quality to some extent. We then select the pose estimation that produces the best compactness value (lowest) and apply it to images within that category. This pipeline helps ensure more accurate pose estimation and improves the overall quality of the 3D reconstruction process. This pipeline is shown in Figure. 7 (a). Generated 3D shapes for the category ”Binderclip” is shown in Figure. 8.
Figure 8: Visualization of Data Preprocessing Results. The first image represents the original figure, followed by the point cloud generated from initial pose estimation. The third image depicts the point cloud reconstruction through refined pose estimation, while the final image showcases the dense point cloud estimation.
3D point cloud Reconstruction: Based on the estimated pose information for each image, we perform multi-view mesh reconstruction, which can then be used for subsequent 3D point cloud-based anomaly detection methods. We following the pipeline established by Neuralangelo [61] to reconstruct the 3D mesh. This is realized by a network that takes a feature considering the 3D point position as input and predicts the surface information. The surface is represented by the zero-level set of the Signed Distance Function (SDF): $S=\{p\in\mathbb{R}^{3}|\text{sdf}_{\theta}(p)=0\}.$ Following Neuralangelo [61], we compute opaque density function $\rho(t)$ from $S$. To make the depth supervision differential, we could derive the rendered depth as in [68]:
[…]
where $t$ is the z value from the sampled point to the camera origin and $N_{S}$ is number of sampled points along each ray. Following [69, 70], give a point $p_{t}$ on the ray of a target frame, a network takes the position and view direction as input to predict the color information. Based on the volume rendering in NeRF [9], the final color is computed as: $\hat{C}=\sum_{j=1}^{N_{S}}T_{j}\alpha_{j}c_{j},$ where $c_{j}$ is the predicted radiance and $N_{V}$ is the number of source views. This structure is shown in Figure. 7 (b).
Discussion: Why it is RGB-Based? The primary reason we do not provide depth images is that the LIDAR sensor cannot accurately predict point positions when the object is close to the camera. However, to capture detailed information about the objects, it is necessary to position the camera near them. This is especially crucial in our scenario, where most of the objects are small. Why this dataset is essential ? Despite the advancements made with current anomaly detection datasets, most of them focus on simulation environments or real-life datasets with ideal conditions, which differ significantly from noisy real-world settings. To bridge this gap, we propose a dataset that accounts for multiple types of noise, including diverse object positions, varying lighting conditions, noisy robot movements, and imperfect camera calibration. We also offer a potential pipeline to address these challenges. However, despite our efforts, this dataset still presents unresolved challenges.
5 Experiments
Problem Definition: For each object, we provide $N$ images for the normal case and $M$ images for cases with defects. The model is trained using the $N$ images or the 3D information estimated from these images, such as point clouds. During testing, the model is used to determine whether an image from the $M$ images contains defects and to generate a mask identifying the defects. In the RAD case, we consider four different types of defects, including both texture-based and geometry-based defects. Challenge: (a) The training and testing images are not aligned, in terms of object pose. (b) The training and testing images are subjected to different lighting conditions. (c) The 3D poses of the training and testing images are not accurate due to camera calibration noise.
Figure 9: Visualization of Pixel-Wise Anomaly Detection Baselines. The heatmap visualization illustrates the ground truth and inference results for two anomalous objects, Bowl and Spraybottle.
Table 3: Quantitative comparison. This table shows the performance of various models on the RAD dataset, evaluated in terms of Pixelwise-ROCAUC and Imagewise-ROCAUC. P-ROC refers to Pixelwise-ROCAUC. I-ROC refers to Imagewise-ROCAUC.
Model Cate. blinder. bowl. box can charger cup1. cup2.
FastFlow [71] P-ROC 65.33 % 67.90 % 74.60 % 66.30 % 99.50 % 65.60 % 71.15 %
I-ROC 89.60 % 78.70 % 97.80 % 88.10 % 81.10 % 86.60 % 71.43 %
STPM [72] P-ROC 67.83 % 73.50 % 78.50% 67.60% 55.80% 74.90 % 74.15 %
I-ROC 75.87 % 70.90 % 87.10 % 72.40% 88.50% 88.30 % 71.15 %
CFA [73] P-ROC 68.41 % 70.29 % 76.78% 67.58% 79.01% 75.14 % 72.53 %
I-ROC 88.60 % 79.44 % 99.67 % 89.43% 92.62% 93.87% 71.30 %
cflow [29] P-ROC 68.45 % 73.70 % 79.26% 67.78% 71.18% 75.72 % 74.52 %
I-ROC 85.38 % 83.76 % 99.89% 92.41% 91.80 % 93.87% 73.33%
CDO [74] P-ROC 96.42 % 60.33 % 84.67 % 87.77 % 97.15 % 81.85 % 60.63 %
I-ROC 58.63 % 64.89 % 72.06 % 50.53 % 66.96 % 43.82 % 84.10 %
UniAD [21] P-ROC 98.75 % 93.88 % 98.58 % 98.69 % 97.88 % 95.56 % 92.28 %
I-ROC 66.30 % 43.09 % 71.02 % 59.53 % 64.62 % 59.71 % 51.81 %
FAVAE [75] P-ROC 57.20 % 57.30 % 35.50 % 83.90 % 59.60 % 45.70 % 62.60 %
I-ROC 49.10 % 53.50 % 52.10 % 33.50 % 53.40 % 60.70 % 29.50 %
PAD [8] P-ROC 90.50 % 75.80 % 90.30 % 98.10 % 97.30 % 96.10 % 79.20 %
I-ROC 48.0 % 37.00 % 62.00 % 46.30 % 51.50 % 48.50 % 54.40 %
Splat [10] P-ROC 98.30 % 96.80 % 97.90 % 98.20 % 99.70 % 98.00 % 97.45 %
I-ROC 52.07 % 80.60 % 55.50 % 37.70 % 57.40 % 81.20 % 48.10 %
Model Cate. glue. phone. rubber. spoon. spray. tennis. Aver.
FastFlow [71] P-ROC 75.60 % 74.43 % 73.40 % 72.00 % 74.00 % 71.00 % 73.14 %
I-ROC 77.85 % 94.78 % 97.10 % 82.30 % 85.50 % 98.00 % 86.83 %
STPM [72] P-ROC 76.75 % 82.35 % 73.80 % 72.60 % 74.60 % 72.00 % 72.64 %
I-ROC 69.75% 85.85 % 77.00 % 79.90 % 82.80 % 98.00 % 80.58 %
CFA [73] P-ROC 76.40 % 74.33 % 71.71 % 74.96 % 73.63 % 84.39 % 74.24 %
I-ROC 80.46% 93.60 % 78.03 % 86.53 % 84.28 % 97.00 % 87.29 %
cflow [29] P-ROC 77.00 % 76.35 % 74.12 % 77.18 % 76.79 % 79.07 % 74.70 %
I-ROC 81.06 % 95.96 % 98.50 % 89.85 % 83.02 % 98.00 % 89.76 %
CDO [74] P-ROC 95.93 % 87.86 % 90.47 % 97.21 % 83.12 % 97.50 % 86.22 %
I-ROC 75.57 % 63.65 % 46.14 % 70.76 % 52.98 % 55.27 % 61.95 %
UniAD [21] P-ROC 97.97 % 97.57 % 98.16 % 99.31 % 98.01 % 99.35 % 97.38 %
I-ROC 56.37 % 71.88 % 65.50 % 75.92 % 57.90 % 68.53 % 62.48 %
FAVAE [75] P-ROC 60.10 % 24.60 % 90.40 % 59.40 % 55.40 % 87.50 % 60.20 %
I-ROC 66.10 % 47.20 % 41.40 % 71.40 % 44.10 % 41.10 % 49.50 %
PAD [8] P-ROC 95.00 % 76.60 % 90.70 % 94.60 % 74.40 % 98.40 % 89.00%
I-ROC 62.60 % 53.40 % 33.00 % 37.30 % 51.60 % 37.60 % 47.94%
Splat [10] P-ROC 99.05 % 97.48 % 96.80 % 99.20 % 92.40 % 98.80 % 97.70 %
I-ROC 42.25% 32.65 % 26.90 % 70.90 % 60.20 % 91.60 % 56.70 %
5.1 Image-Based Baselines
We introduce a dataset comprising multi-view images and evaluate its fidelity using 2D feature embedding-based anomaly detection methodologies, specifically FastFlow [71], STPM [72], CFA [73], cflow [29] and CDO [74]. 2D reconstruction-based anomaly detection methodologies, specifically FAVAE [75] and UniAD [21]. These methods leverage neural networks to learn representations conducive to accurately reconstructing normal samples, with anomalies detected based on deviations in the reconstruction quality. Our findings are summarized in Table. 3. Additionally, we investigate the effectiveness of PAD [8] and Splat [10], a 3D multi-view reconstruction-based algorithm. Notably, both 2D and 3D reconstruction-based methods exhibit limitations. The discrepancy is primarily attributed to misalignment between the training and testing sets, compounded by variations in lighting conditions that impact image quality. Our visualization is also consistent with our findings shown in Figure. 9.
Figure 10: Visualization of Point-Wise Anomaly Detection Baselines. The heatmap visualization illustrates the ground truth and inference results for two anomalous objects, Bowl and Spraybottle.
5.2 Point Cloud-Based Baselines
Additionally, we offer reconstructed point clouds derived from the multi-view RGB inputs. Our experimentation extends to include results obtained through the utilization of the 3D point-cloud-based algorithm BFT [76] shown in Figure. 10. Notably, it is imperative to acknowledge the inherent instability of current multi-view-based surface reconstruction algorithms. These algorithms are sensitive to variations in both lighting conditions and pose estimation, leading to inconsistent point cloud reconstructions across different categories. This inconsistency directly leads to the failure of the 3D point cloud-based anomaly detection algorithm BFT on our dataset. However, the generated point clouds also demonstrate the potential for applying 3D point cloud-based anomaly detection algorithms in multi-view anomaly scenarios, once the pose estimation and 3D shape reconstruction algorithms are properly implemented.
Table 4: Statistics for the RAD Subset. This subset is designed to evaluate the anomaly detection performance of different models across various defect types. For each object in the subset, an equal number of RGB frames is provided for all four defect categories.
Category/Defect Missing Stained Scratched Squeezed Normal Total
bowl_upright 49 49 49 49 220 367
cup1_upright 56 56 56 56 176 344
rubberduck 33 33 33 33 272 371
spraybottle 42 42 42 42 168 328
Table 5: Quantitative comparison. This table shows the performance of various models on the subset of RAD dataset, evaluated in terms of Pixelwise-ROCAUC and Imagewise-ROCAUC.
Model Model Defect bowl cup1 rubber. spray.
FastFlow [71] Pixel ROCAUC Missing 85.00% 89.00% 91.00% 83.00%
Image ROCAUC Missing 87.00% 88.00% 97.00% 94.00%
Pixel Recall Missing 58.00% 83.00% 81.00% 82.00%
FastFlow [71] Pixel ROCAUC Scratched 55.00% 62.00% 65.00% 57.00%
Image ROCAUC Scratched 62.00% 54.00% 89.00% 77.00%
Pixel Recall Scratched 54.00% 52.00% 64.00% 55.00%
FastFlow [71] Pixel ROCAUC Stained 77.00% 80.00% 82.00% 81.00%
Image ROCAUC Stained 96.00% 83.00% 93.00% 91.00%
Pixel Recall Stained 77.00% 78.00% 77.00% 93.00%
FastFlow [71] Pixel ROCAUC Squeezed 59.00% 55.00% 56.00% 65.00%
Image ROCAUC Squeezed 70.00% 63.00% 89.00% 77.00%
Pixel Recall Squeezed 61.00% 53.00% 54.00% 65.00%
STPM [72] Pixel ROCAUC Missing 86.00% 87.00% 87.00% 87.00%
Image ROCAUC Missing 88.00% 89.00% 86.00% 97.00%
Pixel Recall Missing 84.00% 87.00% 88.00% 85.00%
STPM [72] Pixel ROCAUC Scratched 80.00% 81.00% 80.00% 81.00%
Image ROCAUC Scratched 83.00% 80.00% 86.00% 91.00%
Pixel Recall Scratched 80.00% 83.00% 80.00% 84.00%
STPM [72] Pixel ROCAUC Stained 61.00% 63.00% 59.00% 63.00%
Image ROCAUC Stained 66.00% 65.00% 60.00% 65.00%
Pixel Recall Stained 61.00% 64.00% 59.00% 62.00%
STPM [72] Pixel ROCAUC Squeezed 64.00% 65.00% 66.00% 68.00%
Image ROCAUC Squeezed 62.00% 66.00% 68.00% 74.00%
Pixel Recall Squeezed 67.00% 68.00% 67.00% 66.00%
5.3 Ablation Study
In the previous section, we tested the anomaly detection performance of various algorithms across different categories of our dataset. Considering that our dataset also contains objects with various defect types, it is valuable to assess the model’s performance across these defects. Due to the differing materials and shapes of the objects, the number of defect images varies between objects. To address this, we extracted a subset from the original dataset to create a balanced set, ensuring an equal number of frames for each defect type. The statistics for this ”new” dataset are provided in Table. 4. Given time and computational limitations, we applied two baselines to this balanced dataset. The quantitative results, shown in Table. 5, indicate that ”scratched” and ”squeezed” defects pose greater challenges for the baselines compared to ”missing” and ”stained” defects.
6 Disscusion
Conclusion: In this study, we introduce a novel anomaly detection dataset specifically designed to evaluate the robustness of existing anomaly detection algorithms when applied to real-world datasets captured by robots in the presence of environmental noise. Our dataset is characterized by multiple viewpoints and various defects for each object. Based on this dataset, we present a comprehensive preprocessing methodology aimed at facilitating the effective utilization of both 2D and 3D-based anomaly detection algorithms. Additionally, we assess the performance of different algorithms across various categories and defects within our dataset. Limitation: The categories offered within our dataset remain constrained, and the background against which data collection occurred is relatively uncomplicated. Moving forward, our future research endeavors will focus on expanding the range of categories provided.

### v2

@@
 To address this gap, we introduce the Realistic Anomaly Detection (RAD) dataset, the first multi-view RGB-based anomaly detection dataset specifically collected using a real robot arm, providing unique and realistic data scenarios.
-RAD comprises 4765 images across 13 categories and 4 defect types, collected from more than 50 viewpoints, providing a comprehensive and realistic benchmark.
+RAD comprises 4287 images across 13 categories and 4 defect types, collected from more than 60 viewpoints, providing a comprehensive and realistic benchmark.
 This multi-viewpoint setup mirrors real-world conditions where anomalies may not be detectable from every perspective.
-Moreover, by sampling varying numbers of views, the algorithm’s performance can be comprehensively evaluated across different viewpoints.
+Moreover, by sampling varying numbers of views, the baselines’ performance can be comprehensively evaluated across different viewpoints.
 This approach enhances the thoroughness of performance assessment and helps improve the algorithm’s robustness.
-Besides, to support 3D multi-view reconstruction algorithms, we propose a data augmentation method to improve the accuracy of pose estimation and facilitate the reconstruction of 3D point clouds.
-We systematically evaluate state-of-the-art RGB-based and point cloud-based models using RAD, identifying limitations and future research directions.
+Besides, we propose a preprocessed method to support 3D multi-view-based anomaly detection algorithms by improving pose estimation accuracy.
+We systematically evaluate state-of-the-art RGB-based models using RAD, identifying limitations and future research directions.
 The code and dataset could found https://github.com/kaichen-z/PAD.
+1.
+proof-reading.
+2.
+re-arrange code and dataset.
+3.
+re-run experiments.
+(for new dataset) 4.
+make sure the final version.
 keywords:
@@
 Consequently, implementing reliable anomaly detection in real-world robotic systems presents several challenges.
-One major challenge stems from the presence of noise in robotic observations, which can be attributed to various external and internal factors.
+One major challenge stems from the presence of noise in robotic observations, which can be attributed to various external and internal factors [1, 2, 3].
 External factors, such as fluctuating lighting conditions, varying object poses, and inconsistent backgrounds, can introduce significant noise into the data.
@@
 However, obtaining a complete point cloud of an object in real-life conditions is challenging.
-In practical industrial scenarios, objects are not typically suspended in mid-air and are often placed randomly, complicating the acquisition of comprehensive point cloud data.
+In practical industrial scenarios, objects are not typically suspended in mid-air and are often placed randomly, complicating comprehensive point cloud data acquisition.
 Instead of directly providing point cloud information, PAD [11] offers multi-view RGB data, which can be considered a 3D representation akin to Neural Radiance Fields (NeRF) [12].
@@
 These limitations highlight the need for a more robust and comprehensive dataset that can accurately reflect the complexities and variabilities of real-world industrial environments.
+Furthermore, Real-IAD [13] introduced a large-scale, real-world, multi-view industrial anomaly detection dataset.
+However, this dataset provides only five views per object, while in real-world scenarios, objects can appear in a range of random poses, making five fixed observations insufficient to fully capture the variability encountered in practice.
 Based on the analysis above, we introduce RAD, a multi-view RGB-based anomaly detection dataset designed to test the robustness and scalability of existing anomaly detection algorithms on real-world data.
-Our dataset includes 13 different objects and captures over 5000 images from multiple viewpoints (around 50s), providing a comprehensive testbed for evaluating algorithm performance.
+Our dataset includes 13 different objects and captures over 4287 images from multiple viewpoints (around 60s), providing a comprehensive testbed for evaluating algorithm performance.
 RAD records four distinct types of defects, encompassing both texture-based and geometry-based anomalies for each object.
@@
 The main contributions of this dataset are summarized as follow:
-  • We introduce the first real-life multi-view anomaly detection dataset with robotic observations, featuring over 10 object categories and both texture-based and geometry-based defects.
+  • We introduce the first real-life multi-view anomaly detection dataset with robotic observations, featuring 13 object categories and both texture-based and geometry-based defects.
 The dataset effectively captures the noisy environments experienced by real robots, bridging the gap between research and real-world applications.
@@
 Surface represents the objects that are transparent or specular.
-”R.” refers to the dataset with a real-life environment.
-”S.” refers to the dataset with simulation environment.
+“R.”refers to the dataset with a real-life environment.
+“S.”refers to the dataset with simulation environment.
 Datasets Representation #Class # Normal #Abnormal #Shoots
@@
 Real-IAD [13] (R.) RGB 30 99,721 51,329 5
-RAD (Ours) (R.) RGB 13 2,535 2,230 68
+RAD (Ours) (R.) RGB 13 1,224 3,063 68
 Datasets View Color Surface Train. Pose Test. Pose
@@
 Each angle’s coordinates are precisely recorded to maintain uniformity in data capture.
-Nevertheless, owing to the mentioned noises in the introduction, we exclusively offer RGB data and between 40 to 68 views for each category, shown in Figure.
+Nevertheless, owing to the mentioned noises in the introduction, we exclusively offer RGB data and 68 views for each category, shown in Figure.
 2 and Figure.
@@
 Details of data annotation
-To annotate defects, our pipeline comprises two main parts.
-First, we annotate the defect regions in the abnormal image.
-Then, we compare this annotation with the normal mask derived from the corresponding normal image.
+Our pipeline comprises two main parts for annotating defects.
+First, we annotate the object regions in the normal image.
+Then, we compare this annotation with the abnormal image.
 By combining these annotations, we generate a final defect mask.
@@
 Category Type Attribute Miss. St. Sc. Sq. Normal Total
-Binderclip 2 Single 57 57 57 0 171 342
-Bowl 1 Single 55 55 49 55 220 434
-Box 1 Single 44 68 68 0 204 384
-Can 1 Multi./Spec. 41 41 0 41 164 287
-Rubberduck 1 Multi. 68 59 58 68 272 525
-Spraybottle 1 Single/Spec. 37 42 38 42 168 327
-Cup1 1 Single 44 42 44 44 176 350
-Gluebottle 2 Multi./Spec. 32 2 1 44 176 255
-Charger 1 Multi./Spec. 28 28 5 0 84 145
-Cup2 2 Single 67 28 40 67 268 470
-Phonecase 2 Single 68 68 68 68 272 544
-Spoon 1 Single 42 42 25 42 168 319
-Tennisball 1 Multi. 63 64 64 0 192 383
-Labor and time-consuming
+Binderclip 2 Single 136 136 136 0 136 544
+Bowl 1 Single 68 68 60 68 68 332
+Box 1 Single 45 68 68 0 68 249
+Can 1 Multi./Spec. 41 68 4 68 68 249
+Rubberduck 1 Multi. 58 60 68 68 68 322
+Spraybottle 1 Single/Spec. 52 62 68 68 68 318
+Cup1 1 Single 68 66 61 68 68 331
+Gluebottle 2 Multi./Spec. 51 23 24 68 136 302
+Charger 1 Multi./Spec. 66 66 5 0 68 205
+Cup2 3 Single 68 29 66 68 204 435
+Phonecase 2 Single 68 68 68 68 136 408
+Spoon 1 Single 68 66 51 68 68 321
+Tennisball 1 Multi. 67 68 68 0 68 271
+Labor and time-consumption
 The use of a robotic arm for data acquisition represents a significant enhancement in the efficiency and scalability of industrial anomaly detection.
@@
 This meticulous process of annotating images from multiple perspectives ensures a comprehensive representation of potential anomalies but took approximately one months to complete for the dataset.
-Figure 6: Dataset Metrics. (a) illustrates the pixel-wise ratio within each defect across various categories. (b) shows the total number ratio of each defect in RAD.
+Figure 6: Dataset Metrics.
+(a) illustrates the pixel-wise ratio within each defect across various categories.
+(b) shows the total number ratio of each defect in RAD.
+”Mi.” refers to Missing.
+”No.” refers to Normal.
+”Sq.” refers to Squeezed.
+”Sc.” refers to scratched.
+”St.” refers to ”Stained”.
 3.2 Dataset Statistics
@@
 The categories in RAD dataset exhibit significant diversity in shapes and appearances, providing extensive semantic and geometrical exploration spaces that facilitate a wide range of research objectives.
-In addition, the data set also includes metadata (camera pose) for each image, which are essential for nuanced analyzes and model training enhancements.
-3.3 Ground-truth Annotation
-We deliver precise and realistic annotations for each anomaly instance within the test dataset.
-The annotations are represented as two-dimensional regions, ensuring that the pixel positions in the original image correspond exactly to those in the annotated image.
-The normal regions of the objects are delineated in white, whereas the anomalous regions are indicated in different colors.
+In addition, the data set also includes metadata (camera pose) for each image, which is essential for nuanced analyses and model training enhancements.
 Figure 7: Pipeline of Data Preprocessing. The data preprocessing involves two steps: the first step is pose estimation, followed by 3D reconstruction in the second step.
@@
 Table 3: Quantitative comparison.
-This table shows the performance of various models on the RAD dataset, evaluated in terms of Pixelwise-ROCAUC and Imagewise-ROCAUC.
+This table shows the performance of various models on the RAD dataset with masking the background.
+The dataset is evaluated in terms of Pixelwise-ROCAUC and Imagewise-ROCAUC.
 P-ROC refers to Pixelwise-ROCAUC.
 I-ROC refers to Imagewise-ROCAUC.
-Model Cate. blinder. bowl. box can charger cup1. cup2.
-FastFlow [71] P-ROC 65.33 % 67.90 % 74.60 % 66.30 % 99.50 % 65.60 % 71.15 %
-I-ROC 89.60 % 78.70 % 97.80 % 88.10 % 81.10 % 86.60 % 71.43 %
-STPM [72] P-ROC 67.83 % 73.50 % 78.50% 67.60% 55.80% 74.90 % 74.15 %
-I-ROC 75.87 % 70.90 % 87.10 % 72.40% 88.50% 88.30 % 71.15 %
-CFA [73] P-ROC 68.41 % 70.29 % 76.78% 67.58% 79.01% 75.14 % 72.53 %
-I-ROC 88.60 % 79.44 % 99.67 % 89.43% 92.62% 93.87% 71.30 %
-cflow [29] P-ROC 68.45 % 73.70 % 79.26% 67.78% 71.18% 75.72 % 74.52 %
-I-ROC 85.38 % 83.76 % 99.89% 92.41% 91.80 % 93.87% 73.33%
-CDO [74] P-ROC 96.42 % 60.33 % 84.67 % 87.77 % 97.15 % 81.85 % 60.63 %
-I-ROC 58.63 % 64.89 % 72.06 % 50.53 % 66.96 % 43.82 % 84.10 %
-UniAD [21] P-ROC 98.75 % 93.88 % 98.58 % 98.69 % 97.88 % 95.56 % 92.28 %
-I-ROC 66.30 % 43.09 % 71.02 % 59.53 % 64.62 % 59.71 % 51.81 %
-FAVAE [75] P-ROC 57.20 % 57.30 % 35.50 % 83.90 % 59.60 % 45.70 % 62.60 %
-I-ROC 49.10 % 53.50 % 52.10 % 33.50 % 53.40 % 60.70 % 29.50 %
+Model Metric binder. bowl. box can charger cup1. cup2.
+FastFlow [74] P-ROC 94.83% 71.70% 96.60% 90.40% 95.60% 74.10% 92.45%
+I-ROC 69.60% 67.00% 60.30% 49.60% 37.30% 47.50% 76.65%
+STPM [75] P-ROC 94.80% 91.90% 94.30% 95.30% 95.40% 93.80% 94.48%
+I-ROC 69.97% 75.90% 79.50% 68.90% 63.50% 65.50% 69.68%
+CFA [76] P-ROC 99.35% 90.88% 98.99% 97.71% 99.57% 94.61% 95.12%
+I-ROC 90.58% 82.53% 86.74% 92.70% 89.54% 80.52% 73.09%
+cflow [32] P-ROC 99.66% 96.68% 99.66% 99.02% 99.85% 97.97% 97.96%
+I-ROC 88.18% 94.66% 94.17% 95.52% 85.56% 84.45% 77.49%
+Gau P-ROC 99.00% 96.00% 98.20% 98.80% 99.30% 97.30% 96.53%
+I-ROC 21.20% 22.20% 11.80% 27.50% 8.00% 21.20% 19.35%
+UniAD [24] P-ROC 99.21% 95.70% 99.32% 98.90% 99.64% 97.56% 97.29%
+I-ROC 65.74% 63.89% 67.71% 67.28% 51.74% 57.67% 65.10%
+FAVAE [77] P-ROC 99.17% 95.70% 98.20% 99.20% 98.90% 98.30% 97.23%
+I-ROC 57.67% 58.90% 51.90% 55.90% 54.00% 41.30% 54.55%
+CDO [78] P-ROC 99.55% 95.66% 99.50% 98.47% 99.70% 97.43% 97.85%
+I-ROC 80.34% 87.42% 91.10% 86.56% 84.75% 77.99% 79.07%
+Model Metric glue. phone. rubber. spoon. spray. tennis. Aver.
+FastFlow [74] P-ROC 94.20% 96.83% 92.50% 97.00% 93.90% 99.20% 91.49%
+I-ROC 73.85% 91.13% 44.40% 79.50% 80.70% 99.90% 67.49%
+STPM [75] P-ROC 93.70% 93.40% 92.20% 90.20% 91.80% 94.50% 93.52%
+I-ROC 70.15% 79.03% 78.90% 84.70% 74.90% 96.80% 75.19%
+CFA [76] P-ROC 98.72% 97.29% 95.42% 99.00% 95.18% 99.28% 97.01%
+I-ROC 83.21% 81.86% 85.48% 86.83% 88.47% 99.23% 86.21%
+cflow [32] P-ROC 99.39% 98.92% 97.89% 99.69% 97.83% 99.66% 98.78%
+I-ROC 75.87% 99.10% 91.34% 95.83% 83.66% 100.00% 89.68%
+Gau P-ROC 98.95% 97.33% 96.70% 99.30% 94.20% 98.80% 97.72%
+I-ROC 20.80% 21.20% 21.50% 30.00% 20.50% 15.70% 20.07%
+UniAD [24] P-ROC 98.87% 98.54% 97.85% 99.31% 97.87% 99.49% 98.43%
+I-ROC 56.99% 67.24% 64.22% 81.07% 58.83% 66.16% 64.13%
+FAVAE [77] P-ROC 99.10% 97.35% 97.80% 99.50% 97.00% 99.00% 98.19%
+I-ROC 57.35% 42.05% 56.50% 73.20% 59.40% 54.60% 55.18%
+CDO [78] P-ROC 99.11% 98.62% 97.75% 99.23% 96.65% 99.62% 98.40%
+I-ROC 75.52% 92.81% 86.26% 89.24% 65.60% 99.23% 84.30%
+Table 4: Quantitative comparison.
+This table shows the performance of various models on the RAD dataset without masking the background.
+The dataset is evaluated in terms of Pixelwise-ROCAUC and Imagewise-ROCAUC.
+P-ROC refers to Pixelwise-ROCAUC.
+I-ROC refers to Imagewise-ROCAUC.
+Model Metric binder. bowl. box can charger cup1. cup2.
+FastFlow [74] P-ROC 93.73% 68.80% 96.70% 93.30% 94.70% 74.40% 76.10%
+I-ROC 51.17% 60.10% 60.50% 48.30% 40.50% 40.70% 61.80%
+STPM [75] P-ROC 82.73% 75.90% 83.90% 82.60% 85.80% 79.50% 79.58%
+I-ROC 59.90% 53.70% 62.70% 55.10% 49.10% 50.80% 61.28%
+CFA [76] P-ROC 98.00% 59.12% 97.84% 94.53% 98.64% 72.70% 74.37%
+I-ROC 61.49% 58.50% 70.47% 72.99% 61.40% 60.08% 64.06%
+cflow [32] P-ROC 98.83% 85.99% 98.63% 96.94% 99.51% 90.94% 90.18%
+I-ROC 68.14% 66.54% 75.75% 77.29% 80.37% 72.24% 65.32%
+Gau P-ROC 85.37% 64.10% 86.30% 89.30% 87.40% 74.80% 74.15%
+I-ROC 2.27% 2.70% 1.80% 4.70% 2.60% 2.30% 2.50%
+UniAD [24] P-ROC 96.93% 91.04% 96.87% 98.79% 98.04% 96.54% 93.43%
+I-ROC 53.59% 52.19% 52.87% 51.93% 53.28% 55.26% 54.13%
+FAVAE [77] P-ROC 67.90% 39.90% 58.30% 91.10% 83.70% 64.90% 59.65%
+I-ROC 66.20% 20.10% 51.30% 39.50% 35.30% 34.60% 24.78%
+CDO [78] P-ROC 96.65% 69.78% 96.55% 90.52% 97.01% 80.77% 82.12%
+I-ROC 60.51% 48.61% 59.85% 41.56% 55.31% 49.68% 64.93%
 PAD [11] P-ROC 90.50 % 75.80 % 90.30 % 98.10 % 97.30 % 96.10 % 79.20 %
-I-ROC 48.0 % 37.00 % 62.00 % 46.30 % 51.50 % 48.50 % 54.40 %
+I-ROC 48.00 % 37.00 % 62.00 % 46.30 % 51.50 % 48.50 % 54.40 %
 Splat [14] P-ROC 98.30 % 96.80 % 97.90 % 98.20 % 99.70 % 98.00 % 97.45 %
 I-ROC 52.07 % 80.60 % 55.50 % 37.70 % 57.40 % 81.20 % 48.10 %
-Model Cate. glue. phone. rubber. spoon. spray. tennis. Aver.
-FastFlow [71] P-ROC 75.60 % 74.43 % 73.40 % 72.00 % 74.00 % 71.00 % 73.14 %
-I-ROC 77.85 % 94.78 % 97.10 % 82.30 % 85.50 % 98.00 % 86.83 %
-STPM [72] P-ROC 76.75 % 82.35 % 73.80 % 72.60 % 74.60 % 72.00 % 72.64 %
-I-ROC 69.75% 85.85 % 77.00 % 79.90 % 82.80 % 98.00 % 80.58 %
-CFA [73] P-ROC 76.40 % 74.33 % 71.71 % 74.96 % 73.63 % 84.39 % 74.24 %
-I-ROC 80.46% 93.60 % 78.03 % 86.53 % 84.28 % 97.00 % 87.29 %
-cflow [29] P-ROC 77.00 % 76.35 % 74.12 % 77.18 % 76.79 % 79.07 % 74.70 %
-I-ROC 81.06 % 95.96 % 98.50 % 89.85 % 83.02 % 98.00 % 89.76 %
-CDO [74] P-ROC 95.93 % 87.86 % 90.47 % 97.21 % 83.12 % 97.50 % 86.22 %
-I-ROC 75.57 % 63.65 % 46.14 % 70.76 % 52.98 % 55.27 % 61.95 %
-UniAD [21] P-ROC 97.97 % 97.57 % 98.16 % 99.31 % 98.01 % 99.35 % 97.38 %
-I-ROC 56.37 % 71.88 % 65.50 % 75.92 % 57.90 % 68.53 % 62.48 %
-FAVAE [75] P-ROC 60.10 % 24.60 % 90.40 % 59.40 % 55.40 % 87.50 % 60.20 %
-I-ROC 66.10 % 47.20 % 41.40 % 71.40 % 44.10 % 41.10 % 49.50 %
+Model Metric glue. phone. rubber. spoon. spray. tennis. Aver.
+FastFlow [74] P-ROC 91.55% 97.00% 94.10% 96.70% 94.90% 98.20% 90.01%
+I-ROC 36.20% 65.48% 49.20% 78.40% 57.30% 78.50% 56.01%
+STPM [75] P-ROC 85.80% 85.88% 84.80% 79.10% 82.60% 85.40% 82.58%
+I-ROC 56.05% 73.23% 50.80% 59.70% 60.10% 71.50% 58.77%
+CFA [76] P-ROC 95.83% 95.67% 92.93% 98.43% 92.89% 98.76% 89.98%
+I-ROC 56.50% 73.77% 62.91% 73.96% 59.31% 83.53% 66.07%
+cflow [32] P-ROC 97.06% 94.02% 95.14% 98.55% 91.55% 98.50% 95.06%
+I-ROC 66.43% 66.09% 57.61% 82.17% 60.37% 79.37% 70.59%
+Gau P-ROC 78.40% 86.60% 85.70% 96.60% 90.50% 93.70% 84.07%
+I-ROC 1.60% 4.65% 5.60% 14.70% 20.60% 4.00% 5.39%
+UniAD [24] P-ROC 97.15% 96.16% 96.43% 97.69% 91.37% 98.86% 96.10%
+I-ROC 50.45% 51.72% 55.16% 50.02% 50.13% 51.29% 52.46%
+FAVAE [77] P-ROC 74.50% 59.58% 88.00% 85.70% 70.40% 92.00% 71.97%
+I-ROC 46.10% 32.28% 33.20% 53.10% 44.80% 49.30% 40.81%
+CDO [78] P-ROC 94.29% 93.36% 92.35% 96.16% 88.36% 97.26% 90.40%
+I-ROC 58.98% 53.60% 52.97% 53.45% 53.98% 66.34% 55.37%
 PAD [11] P-ROC 95.00 % 76.60 % 90.70 % 94.60 % 74.40 % 98.40 % 89.00%
@@
 Our findings are summarized in Table.
-3.
+3 and Table.
+4.
 Additionally, we investigate the effectiveness of PAD [11] and Splat [14], a 3D multi-view reconstruction-based algorithm.
@@
 The statistics for this ”new” dataset are provided in Table.
-4.
+5.
 Given time and computational limitations, we applied two baselines to this balanced dataset.
 The quantitative results, shown in Table.
-5, indicate that ”scratched” and ”squeezed” defects pose greater challenges for the baselines compared to ”missing” and ”stained” defects.
+6, indicate that ”scratched” and ”squeezed” defects pose greater challenges for the baselines compared to ”missing” and ”stained” defects.
 6 Disscusion